Circular No. 7015

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998dy, 1998dz, 1998ea D. Reiss, University of Washington, reports that the Mount Stromlo

Abell Cluster Supernova Search Team (cf. IAUC 6639) has discovered three apparent supernovae on V and R CCD images taken by S. Sabine at the Mount Stromlo 1.27-m telescope (+ Macho Camera):

SN  1998 UT  α_{2000}  δ_{2000}  V  Offset
1998dy Aug. 1 4^{h}29^{m}54^{s}.24 −61^{o}12′43′′.1 19.0  —
1998dz Sept. 3 3 20 58.18 −41 05 22.8 18.5  —
1998ea Sept. 6 4 33 46.59 −61 35 20.4 19.5 5′′.9 W, 1′′.5 N

SNe 1998dy and 1998ea are near Abell 3266. SNe 1998dy and 1998dz have no apparent host galaxy. SN 1998dy is also visible on frames taken on Aug. 22, 28, and Sept. 9, and it is now fading; a subsequent check of previous images shows that it was also present on images obtained on July 14 (V ∼ 19.0), but SN 1998dy was not visible on similar images taken on 1998 Mar. 16 (limiting mag V ∼ 20.5) or Apr. 13 (limiting mag V ∼ 20). SN 1998dz is near Abell 3122 and was confirmed on images taken on Sept. 6, and it appears to be rising in brightness; it was not visible on similar images taken on July 14 (limiting mag V ∼ 20.0) and Aug. 22 (limiting mag V ∼ 21.0). SN 1998ea was confirmed on images obtained on Sept. 9, and it seems to be brightening; it was not visible on similar images taken on Aug. 28 and 31 (limiting mag V ∼ 20.5). Finding charts are available

at http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

NOVA IN M31 E. Halderson, T. Shefler, M. Modjaz, J. Y. King, M. Papenkova, W. D.

Li, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent nova in M31 (NGC 224) during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). The nova was found and confirmed (unfiltered mag 17.8) on unfiltered CCD images taken on Sept. 15.3 and 16.3 UT, respectively. The object is located at α = 0^{h}42^{m}46^{s}.6, δ = +41^{o}14′50′′ (equinox 2000.0; due to a lack of nearby stars, this semiaccurate position was calculated as an offset from the NED position for the nucleus of M31), which is ∼ 26′′ east and 78′′ south of the galaxy nucleus. A KAIT image of M31 taken on Aug. 31.3 showed nothing at the position of the nova (limiting mag ∼ 19.0).

1998 September 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green